### v4

# Non-coherent evolution of closed weakly interacting system leads to equidistribution of probabilities of microstates

Abstract
We introduce a concept of non-coherent evolution of macroscopic quantum systems. We show that for weakly interacting systems such evolution is a Markovian stochastic process. The transition rates between system states, which characterize the process, are determined by Fermi’s golden rule. Such evolution is time-irreversible and leads to the equidistribution of probabilities across every state of the system. Furthermore, we investigate the time dependence of the mean numbers of particlesin single-particle states and find that, under the given assumptions, it is governed by the Boltzmann collision integral. In this theory, non-coherence is the mechanism that transforms time-reversible unitary evolution into time-irreversible stochastic evolution. Thus, we present a possible solution to the famous arrow-of-time problem.
I Introduction
The arrow-of-time problem, as presented in textbooks on statistical mechanics [1, 2] and in specialized books on the nature of time [3, 4], remains one of the most fundamental challenges in modern physics. The issue arises from the time asymmetry in the equations governing the behavior of large ensembles of particles, in contrast to the time-symmetric nature of the equations describing individual particles.
Let us list the well-known time-asymmetric properties of macroscopic systems. Probably the most famous is the second law of thermodynamics, which states that the entropy of an isolated system undergoing spontaneous evolution cannot decrease [2]. Next is the existence of thermodynamic equilibrium—a state toward which an isolated system evolves. This state is characterized by the equidistribution of probabilities among microstates. The state in which all microstates are equally probable is called the Gibbs microcanonical ensemble [5]. Finally, the Boltzmann equation, which describes the time evolution of single-particle distribution functions, is famously irreversible.
The classical approach to deriving the Boltzmann equation is the BBGKY hierarchy [6]. However, it postulates molecular chaos as a model for the two-particle distribution function without fully elucidating its origins. Another interesting attempt to deduce the Boltzmann equation is found in Refs. [7, 8], where it is treated as a mean-field theory; however, only the diffusive, reversible part can be derived in this way. In Ref. [9], the authors take the reverse route – not deriving the concept of probability equidistribution, but rather arguing that statistical physics can be formulated without this assumption.
Other approaches to explaining asymmetry postulate low-entropy initial states of the universe [10, 11], which is not completely satisfactory, as it constitutes an ad hoc hypothesis [12]. In contrast, Ref. [13] proposes that physical laws themselves must be modified to incorporate irreversibility.
In the textbook [14], the second law of thermodynamics is derived from an equation known as the Fermi master equation. However, it is noted that the equation itself lacks a proper derivation. A development of this idea is the Lindblad master equation [15, 16], which describes systems weakly coupled to their environments. These equations have been shown to exhibit irreversible behavior, even when the underlying microscopic dynamics are unitary. This approach has been applied to various models in quantum thermodynamics [17, 18]. A particularly important result from this line of research is the identification of conditions under which systems reach a steady state that mirrors the classical Boltzmann distribution [19, 20].
Modern approaches propose that irreversibility emerges from a time-asymmetric mechanism inherent in quantum mechanics [21]. Typically, decoherence is suggested as this mechanism [22]. Decoherence theory explains measurement in quantum theory as the interaction of a quantum system with its environment, which causes the loss of coherence in the system [23].
Another notable approach is the study of quantum entanglement and its relationship to thermodynamic irreversibility. Some researchers have proposed that the generation and disruption of entanglement within a quantum system may play a key role in the emergence of macroscopic irreversibility [24, 25].
In [26], the authors derive the second law of thermodynamics from a quantum-mechanical analysis of a total system partitioned into an object and its environment. This type of theory has attracted considerable attention in recent years [27, 28, 29]. Additionally, the theory of non-Markovian quantum evolution [30, 31, 32] has garnered significant interest. It focuses on cases where the system’s dynamics are influenced by interactions with the external world.
However, all of the aforementioned theories locate irreversibility within a subsystem of a larger, isolated system – the environment. The total system, being isolated, is still governed by reversible equations, which represents a significant conceptual gap [35]. In the classical Gibbs treatment of subsystems, they are considered part of a larger system in equilibrium, described by the microcanonical ensemble. Reversible equations, however, cannot reach this stable state [1]. From the perspective of the second law of thermodynamics, satisfying the condition of entropy growth for the entire system requires postulating a low-entropy initial state. But this is the solution, which the authors of the cited works dismiss as ad hoc and unsatisfactory.
To address this, in the present paper we use the concept of decoherence but approach the notion of non-coherence from a different perspective than decoherence theory. Rather than focusing on the loss of coherence in a quantum system due to interaction with its environment, our investigation centers on the environment itself. We treat it as inherently non-coherent. We observe that it is the coherent states of matter that are exotic, while non-coherent states occur naturally. Coherent states – such as superconductive and superfluid phases – arise only in specific substances under low temperatures. Coherent light is generated by lasers. In contrast, natural light is typically non-coherent. From this, we infer that non-coherence represents the inherent, natural state of matter.
In Ref. [33], we proposed the concept of coherent and non-coherent transformation to describe the emergence of irreversibility associated with the transmission of quantum particles through an interface. It was suggested that if particles incident on the interface from opposite sides are phase-correlated (coherent regime), their transmission occurs reversibly. In contrast, transmission is irreversible when there is no phase correlation between particles on the two sides of the interface (non-coherent regime), which is the naturally occurring situation. An optical experiment was conducted and successfully confirmed the predictions of this theory [34]. Extending this theory, we introduce the concept of coherence time into the evolution of an isolated quantum system.
The most significant difference between the approach presented here and decoherence theory is that the latter, in the spirit of the original Copenhagen interpretation, divides changes occurring in a quantum system into two categories: normal evolution, described by the Schrödinger equation, and measurement. The latter is then associated with the process of decoherence. Of these two parts, it is measurement that is considered irreversible [22]. According to this theory, decoherence arises due to the interaction of the quantum system with its environment. However, it is not universally accepted that decoherence theory resolves the measurement problem [35]. We treat decoherence differently, drawing on an analogy with wave optics. In optics, non-coherence arises because no wave with a continuous spectrum is entirely monochromatic. As a result, light emitted from different sources cannot form stable interference patterns, and waves emitted from the same source will desynchronize if split and travel along paths whose length difference exceeds the coherence length [36]. In our view, it is not the process of coherence loss that is irreversible, but rather the evolution of a system large enough that different parts of it are no longer mutually coherent.
We aim to comprehend how a system described by quantum theory but lacking coherence evolves. The mathematical procedure used to derive equations for non-coherent light involves averaging the arguments of the complex amplitudes of waves. A parallel between quantum entanglement and the coherence of light was recently observed in Ref. [37]. Our objective is to extend the procedure of averaging the arguments of complex amplitudes to general quantum systems and derive the equation governing non-coherent evolution.
We employ a method in this regard that is very similar to the one used in decoherence theory to eliminate the non-diagonal components of the density matrix [23]. However, in our approach, the non-diagonal terms that vanish do not belong to the matrix characterizing the state of the system, but rather to the matrix describing the transformation of the system.
After performing the averaging over arguments, we obtain the equation of non-coherent evolution. We find that it coincides with the previously mentioned Fermi master equation. The Fermi master equation is equivalent to Fermi’s golden rule combined with the conservation of total probability [14]. This means that by introducing the concept of non-coherent evolution into quantum mechanics, we arrive at equations that already exist. Thus, we do not contradict any well-known or well-tested results of quantum mechanics. Rather, we re-derive and reinterpret them in a way that makes quantum mechanics compatible with statistical mechanics, particularly with the irreversibility found in kinetic equations. In [14], it is shown that entropy growth follows from the Fermi master equation. We will additionally show that both the equidistribution of probabilities and the Boltzmann equation can be readily derived from it.
II Non-coherent transformations
As stated in the introduction, the concept of non-coherent transformation was introduced in Ref. [33] to describe the transmission of quantum particles through an interface. It was proposed to explain the emergence of irreversibility in such cases. As a physical example of the formalism presented below, one can imagine a beam of quantum particles incident on an interface and being scattered in different directions. However, except for one important case, we will not refer to this specific example and will proceed in an abstract manner.
The goal here is to develop and generalize the formalism introduced in Ref. [33] so that it can be applied in the next section. In this section, we will first deduce the main formula for a series of non-coherent transformations and explain the origin of the argument averaging procedure. The following section will address non-coherent evolution, which is a continuous parametric family of non-coherent transformations where the parameter is time.
Let us consider a system that can occupy a set of different microstates indexed by $k$, taking values from $1$ to $N$. The system may be in different microstates with different probabilities. The probability amplitude of the system being in microstate $k$ is denoted by $A_{k}$, and the probability is given by $P_{k}=|A_{k}|^{2}$. The state of the system is characterized by the complete set of probability amplitudes. We will group all these amplitudes $A_{k}$ into a column.
We introduce an index-free notation for further use, as it is more transparent in certain cases. We denote matrices and column vectors using a specific font: columns of amplitudes as $\mathcal{A}$, columns of probabilities as $\mathcal{P}$, and the relation between probabilities and amplitudes is written as $\mathcal{P}=\mathcal{A}^{*}\circ\mathcal{A}$, where the circle denotes the Hadamard product, or element-wise product. This can be written more concisely by introducing the Hadamard square, or element-wise square:
$\mathcal{P}=|\mathcal{A}|^{\circ 2}.$ (1)
The same notation will be applied to transformation matrices later.
Let the system evolve freely, without interactions between microstates. This implies conservation of probabilities—and hence the absolute values of amplitudes—but allows for changes in the phases of the amplitudes corresponding to different microstates. Let us denote by $\Phi$ the matrix that characterizes the phase changes. By definition, it is diagonal:
$\Phi_{kj}=e^{i\varphi_{k}}\delta_{kj}.$ (2)
At some moment, the system that was initially in state $i$ undergoes changes, leading to a new state $f$, characterized by a column of amplitudes $\mathcal{A}_{f}$. The linear transformation between states should be unitary to conserve total probability, and we denote the unitary transformation matrix as $\mathcal{U}$. Thus, the new state of the system $\mathcal{A}_{f}$ is given by
$\mathcal{A}_{f}=\mathcal{U}\Phi\mathcal{A}_{i}$
$\mathcal{P}_{f}=|\mathcal{U}\Phi\mathcal{A}_{i}|^{\circ 2}.$ (3)
Let us rewrite the last formula component-wise. We denote $k_{i}$ indexes of microstates of the initial state and $k_{f}$ of the final state
$P_{k_{f}}=|A_{k_{f}}|^{2}=\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}+\sum_{k^{\prime}_{i}\neq k_{i}}U^{*}_{k_{f}k^{\prime}_{i}}U_{k_{f}k_{i}}e^{i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})}A^{*}_{k^{\prime}_{i}}A_{k_{i}}$ (4)
We impose the non-coherence condition on the transformation, signifying that we lack knowledge of the arguments $\varphi$ of the amplitudes in Eq. (4); only their absolute values are known. To ascertain how the probabilities transform, we average over all possible values of the arguments:
$P_{k_{f}}=\frac{1}{(2\pi)^{N}}\idotsint\limits_{0}^{2\pi}\prod_{k_{i}}d\varphi_{k_{i}}\left[\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}+\sum_{k^{\prime}_{i}\neq k_{i}}U^{*}_{k_{f}k^{\prime}_{i}}U_{k_{f}k_{i}}e^{i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})}A^{*}_{k^{\prime}_{i}}A_{k_{i}}\right].$ (5)
[…]
$P_{k_{f}}=\sum_{\beta}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}$
$\mathcal{P}_{f}=|\mathcal{U}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}$
$\mathcal{P}_{f}=\mathcal{T}\mathcal{P}_{i}.$ (6)
Here we introduced the matrix $\mathcal{T}=|\mathcal{U}|^{\circ 2}$, the matrix of non-coherent transformation. What are the physical consequences of having the transformation law (6) instead of (3)?
First of all, we observe that since we average over the arguments, the phases of the initial amplitudes $A_{k_{i}}$ play no role in the expression for the probabilities of the final state (6). Therefore, we do not need to consider the initial column of amplitudes $\mathcal{A}$ and can work directly with the column of probabilities $\mathcal{P}$.
Second, it follows from the properties of unitary matrices that the matrix $\mathcal{T}$ is bistochastic, meaning its elements are non-negative real numbers, and the sum of the elements in each row and each column equals one. Since the sum of the elements in each column is one, the values $T_{k_{f}k_{i}}$ can be interpreted as the probabilities of transition from state $k_{i}$ to state $k_{f}$.
The third, and most important, consequence is the irreversibility of non-coherent transformation. Bistochastic matrices generally lack inverses within the same class of matrices (except for permutation matrices, which consist solely of ones and zeroes). The physical consequence of this is that transformations described by such matrices—other than permutations—are irreversible. A detailed discussion of this is given in Ref. [33].
So, if we are allowed to average over arguments, we obtain irreversible behavior of the system under consideration. But what gives us the right to perform such averaging? To answer this question, we need to consider concrete physical examples.
Figure 1: Two beams of light are incident on the interface and are transformed either coherently or non-coherently, depending on the conditions. The interface can be, for example, between air and optical glass (depicted as light blue). The gray rectangles represent mirrors. In sub-pictures (a) and (b), two beams are emitted by the same source and separated by the interface. The initial state of the system is defined by the amplitudes after separation. The column of initial amplitudes is denoted as $\mathcal{A}_{i}$, and the column of initial intensities as $\mathcal{P}_{i}$.
If the difference between the optical paths of the two split beams is smaller than the coherence length of the light, $\Delta l<l_{\text{coh}}$, a coherent transformation occurs according to formula (3), as depicted in sub-picture (a). This transformation is reversible, and with appropriate phase differences between the beams, they can merge back into a single beam, meaning the splitting transformation has been reverted. If the difference between the optical paths exceeds the coherence length of the light, $\Delta l>l_{\text{coh}}$, a non-coherent transformation occurs according to formula (6). The resulting intensities no longer depend on the phase difference and are strictly greater than zero, which means the split beams cannot merge back into a single beam. In sub-picture (c), two beams from different sources meet at the interface. They cannot form a stable interference pattern, which means they transform irreversibly, in accordance with formula (6).
In Ref. [34], an optical experiment is conducted to demonstrate the irreversibility of non-coherent transmission of light through the interface. First, a beam of light is split into two parts by the interface between air and optical glass. Then, after reflecting off mirrors, the two parts meet again at the interface (Fig. (1)a, b). If the difference in the optical path lengths exceeds the coherence length, $\Delta l>l_{\text{coh}}$, the resulting intensities of the beams can be calculated using a product of bistochastic matrices, without considering the phase difference between the beams. In this case, formula (6) should be used. However, if the difference in the optical path is smaller than the coherence length, $\Delta l<l_{\text{coh}}$, the actual phase difference between the two parts of the light must be taken into account, so we calculate using formula (3). Importantly, in the latter case, if the phase difference between the two beams differs by $2\pi n$, the beams reunite into one at the interface—demonstrating the reversibility of the process.
Let us abstract away the details of the described experiment and formulate the mathematical essence. The formulas we provide in the following few paragraphs are well-known in the contexts of classical [36] and quantum optics [Knight], so we will not provide extensive detail. However, we aim to rewrite them in the specific context and notation of this manuscript for further use.
Also, it will be more convenient to refer to the coherence time $t_{\text{coh}}$ rather than the coherence length $l_{\text{coh}}$. The two quantities are related by the simple formula $l_{\text{coh}}=vt_{\text{coh}}$, where $v$ is the phase velocity of the wave.
If two monochromatic waves with amplitudes $A_{1}$ and $A_{2}$ overlap each other, the resulting intensity is
$I_{\text{res}}=|A_{1}|^{2}+|A_{2}|^{2}+2|A_{1}||A_{2}|\cos\phi,$ (7)
where $\phi$ is the difference in phase between the two waves. The term with the cosine is the interference term, which we denote as Int. However, we know that in practice, no wave is completely monochromatic. This means that, in frequency space, it is not a delta function, but rather a collection of many frequencies, even if they are concentrated near some given value. The amplitude at a given time is then given by the expression
[…]
Considering this, the interference term is given by
[…]
This formula is valid if we calculate the average intensity over a sufficiently large time interval. On smaller time scales, the intensity may fluctuate above or below this value, since the formula relies on the orthogonality of exponential functions with different frequencies – something that is strictly true only for integration over an infinite time. However, if the averaging time interval is much greater than the inverse of the frequency difference, the integral becomes close to zero.
In the experiment described above and depicted in Fig. (1a, 1b), the wave interacts with itself, but taken at a different moment in time. We substitute into the formula (9) the expressions $A_{1}(t)=A(t)$ and $A_{2}(t)=A(t+\Delta t)$, and for these we use the expression (8), so we obtain
[…]
It can be shown that such an integral vanishes when $\Delta t\to\infty$. To gain a better understanding of the behavior of the interference as a function of $\Delta t$, we need to make some additional assumptions. The standard approach is to assume $|A(\omega)|=\mathrm{const}$ within some frequency interval, and zero outside of this interval [36]. We denote the spectral width of the interval as $\omega_{\text{width}}$, and we can approximate the integral (10) as being of order one when $\Delta t<1/\omega_{\text{width}}$, and of order $1/(\Delta t\omega_{\text{width}})$ otherwise. We define the coherence time $t_{\text{coh}}$ as $1/\omega_{\text{width}}$, and rewrite these estimations as
[$&\Delta t<t_{coh},\,\text{Int}=\,1\\ &\Delta t>t_{\text{coh}},\,\text{Int}=\frac{t_{\text{coh}}}{\Delta t}\end{aligned}\right.$ (11)
We conclude that when not fully monochromatic and, generally, not fully correlated waves overlap, the resulting intensity is given by
$I_{res}=|A_{1}|^{2}+|A_{2}|^{2}+2\,\text{Int}\cdot|A_{1}||A_{2}|\cos\phi,$ (12)
where Int is given by expression (11). We see that if $\Delta t\gg t_{coh}$, the resulting intensity is simply the sum of the intensities of the two sources. This case is referred to as the non-coherent regime.
We can explain the non-coherent regime differently. The interaction of each single-frequency component of $A_{1}$ in the formula (9) with the same-frequency component of $A_{2}$ is described by expression (7). However, the phase difference $\phi$ in that formula is a function of frequency. When calculating Int, we integrate $\cos\phi$ over all frequencies and obtain zero.
Another case where averaging over phases is applicable is when particles emitted at different points meet at the interface. If the emissions of these particles are not causally related, they cannot exhibit stable interference. The optical analog of this was studied in Refs. [38, 39], where the authors investigated interference from two different sources. While interference patterns form at any given moment, over a time interval comparable to the decoherence time, the interference pattern drifts: constructive interference may turn into destructive interference and vice versa. Therefore, when calculating the mean intensity over a time interval much greater than the coherence time, we should average over phases.
A mathematical explanation of this may be the following. If we restrict both $A_{1}$ and $A_{2}$ from formula (9) so that they are nonzero only within some finite frequency interval, the formula defines a scalar product in an infinite-dimensional Hilbert space. In $n$-dimensional space, the expected value of the scalar product of two unit vectors is proportional to $n^{-1/2}$ [40]. So as $n\to\infty$, it tends to zero. This means that if we pick two random sources, their interference defined by formula (9) is zero.
Thus, the two distinct mechanisms leading to a lack of coherence in interaction, depicted in Fig. ((1)b, (1) c), yield the same result: the correct outcome is found by averaging over phases.
We have described two mechanisms leading to the switch from coherent transformation to the non-coherent one. The first mechanism is particularly interesting because it allows for direct experimental measurements in both coherent and non-coherent regimes, and the transition between these regimes can be described theoretically by formula (LABEL:interference). However, we suppose that the second mechanism – namely, the interaction of waves whose emissions are not causally related – is the one responsible for the non-coherence of transformations of the states of macroscopic-sized matter. In later sections, we will illustrate this with concrete examples.
Finally, let us generalize our results to an arbitrarily long series of transformations. Let the final set of probabilities be expressed as
$\mathcal{P}_{f}=|\mathcal{U}_{n}\Phi_{n}...\mathcal{U}_{1}\Phi_{1}\mathcal{A}_{i}|^{\circ 2},$ (13)
after $n$ transformations. The case where the time between initial ($i$) and final ($f$) states is much less than the coherence time of the system, means we can disregard the phase shift between transformations and thus we have
$\mathcal{P}_{f}=|\mathcal{U}_{n}...\mathcal{U}_{1}\mathcal{A}_{i}|^{\circ 2}..$ (14)
For the large time intervals between transformations, when the conditions for averaging over arguments are fulfilled, we can write
$\mathcal{P}_{f}=\int d^{N}\varphi_{n}\,...\int d^{N}\varphi_{1}\,|\mathcal{U}_{n}\Phi_{n}...\mathcal{U}_{1}\Phi_{1}\mathcal{A}_{i}|^{\circ 2}.$ (15)
where we denoted
$\int d^{N}\varphi_{k}=\frac{1}{(2\pi)^{N}}\idotsint\limits_{0}^{2\pi}\prod_{k_{j}=1}^{N}d\varphi_{k_{j}}$ (16)
To simplify the expression (15) we want to write it in an index form. For (15) we have
$P_{k_{f}}=\sum_{k_{n},k^{\prime}_{n}=1}^{N}...\sum_{k_{1},k^{\prime}_{1}=1}^{N}\sum_{k_{i},k^{\prime}_{i}=1}^{N}(U_{k_{f},k^{\prime}_{n}}e^{i\varphi_{k^{\prime}_{n}}}...U_{k^{\prime}_{1},k^{\prime}_{i}}e^{-i\varphi_{i^{\prime}}}A_{k^{\prime}_{i}})^{*}(U_{k_{f},k_{n}}e^{i\varphi_{k_{n}}}...U_{k_{1},k_{i}}e^{i\varphi_{i}}A_{k_{i}}).$ (17)
We perform summation over each set of microstate indeces $k_{j}$ for every state $j$ from $1$ to $n$.
When we do averaging over phases, couples of exponents $\exp i(\varphi_{k}-\varphi_{k^{\prime}})$ turns to delta symbols $\delta_{k^{\prime},k}$, so after averaging we get
$P_{k_{f}}=\sum_{k_{n},k^{\prime}_{n}=1}^{N}...\sum_{k_{1},k^{\prime}_{1}=1}^{N}\sum_{k_{i},k^{\prime}_{i}=1}^{N}(U_{k_{f},k^{\prime}_{n}}...U_{k^{\prime}_{1},k^{\prime}_{i}}A_{k^{\prime}_{i}})^{*}\delta_{k^{\prime}_{n},k_{n}}...\delta_{k^{\prime}_{i},k_{i}}(U_{k_{f},k_{n}}...U_{k_{1},k_{i}}A_{k_{i}}).$ (18)
We sum over each $k^{\prime}_{j}$ using delta-symbols and rearrange terms
$P_{k_{f}}=\sum_{k_{n}=1}^{N}...\sum_{k_{1}=1}^{N}\sum_{k_{i}=1}^{N}U_{k_{f},k_{n}}^{*}U_{k_{f},k_{n}}...U_{k_{1},k_{i}}^{*}U_{k_{1},k_{i}}A_{k_{i}}^{*}A_{k_{i}}.$ (19)
Which, with indexless notation just gives
$\mathcal{P}_{f}=|\mathcal{U}_{n}|^{\circ 2}...|\mathcal{U}_{1}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}$
$\mathcal{P}_{f}=\mathcal{T}_{n}...\mathcal{T}_{1}\mathcal{P}_{i}.$ (20)
If we compare the derived expression with (14), we find that (20) involves taking Hadamard squares before multiplying the matrices, whereas (14) involves multiplying the matrices first and then taking the Hadamard square.
Now, intuitively, if each $U_{k}$ represents an infinitesimal transformation, or a transformation close to the identity matrix, and there is an infinite number of such transformations, the expression (14) would represent unitary, or coherent evolution, while the expression (20) would represent non-coherent, or stochastic evolution.
In the next section, we will make this idea precise and combine it with basic concepts from the theory of quantum transitions under the influence of perturbations to derive the equation of non-coherent evolution.
Before proceeding to the next section, we want to highlight a serious issue in our reasoning. The Hadamard product is basis-dependent, so we must specify the basis before averaging over phases, as in (6). Physically, this means we must select a specific set of microstates as ”true” and distinguish them from any other state that is a superposition of the chosen microstates. In this manuscript, we consider weakly interacting systems, where the choice is evident: the states of the system can be expressed by the occupation numbers of the one-particle Hamiltonian’s eigenstates. As we will see later, this choice naturally leads us to the correct form of the Boltzmann equations. However, how to choose the set of ”true” microstates in the general case remains unclear. This issue is already recognized in the context of decoherence theory as the superselection problem. Possible approaches to resolving it are discussed in the book [23].
III Non-coherent evolution equation and quantum transitions
As the previous section shows, in the discrete case, the non-coherence condition, $\Delta t\gg t_{\text{coh}}$ or $\Delta E\ll\omega_{\text{width}}$, converts a reversible unitary transformation into an irreversible stochastic transformation. In the continuous case, unitary evolution is described by the exponential of a skew-Hermitian matrix [41]. In the context of quantum theory, we typically express this as $\mathcal{U}(t)=\exp(i\mathcal{H}t)$, where $\mathcal{H}$ is a Hermitian matrix, particularly representing the Hamiltonian of the system. The continuous analog of a bistochastic transformation is a symmetric continuous-time Markov chain (CTMC). The time dependence of transition probabilities is given by $\mathcal{T}(t)=\exp(\mathcal{Q}t)$, where $\mathcal{Q}$ is the transition rate matrix (Q-matrix) [42]. Its non-diagonal elements are positive numbers representing the rates of transition between states. The diagonal elements $Q_{\alpha\alpha}$ represent the total departure rate from the state $\alpha$. They are negative, such that the sums of elements in each column are zero [^1] In probability texts, the state of the system is often represented by a row, and the transformation matrix is placed on the right of it. We use the convention adopted in physical texts: to represent the state as a column and place the transformation matrix to the left of it. This convention results in a difference in definition: in probability textbooks, the sums of elements in each row of a Q-matrix are zero.. A general Q-matrix serves as the continuous analog of a stochastic transformation, while the analog of a bistochastic transformation is a Q-matrix symmetric over the main diagonal, i.e., $Q_{\alpha\beta}=Q_{\beta\alpha}$.
We are going to demonstrate that just as the unitary transformation converts into a bistochastic transformation under the condition of non-coherence, unitary evolution converts into symmetric stochastic evolution. In our consideration, we will extensively use the fact that a closed quantum system has a discrete spectrum. However, we consider a macroscopic system where a continuous spectrum with a well-defined density of states serves as a suitable approximation, which is typical for a closed macroscopic system [1]. We consider interaction in a local area of a closed system so that the emission of interacting particles would typically occur at some distance and be causally unrelated. We will specify this requirement further in Section IV after we have a good understanding of the relation of parameters necessary for the presented theory to be applicable. We also imply the condition of weak interaction, since strongly interacting systems are beyond the applicability of the methods that we use.
We start similarly to a classic simple time-dependent perturbation theory [43] that is aimed to derive Fermi’s golden rule. We seek for a wave function of a many-body system in the form
$\Psi=\sum_{k}A_{k}|k\rangle,$ (21)
where $|k\rangle$ are states of non-interacting system, and $A_{k}$ are probability amplitudes of system to be in state $|k\rangle$. Standartly we get the following equation for $A_{k}$:
$i\hbar\sum_{k}\dot{A}_{k}|k\rangle=\sum_{l}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)\hat{V}A_{l}|l\rangle.$ (22)
We introduce $V_{kl}=\langle k|\hat{V}|l\rangle$ and get the equation for amplitudes
$\dot{A}_{k}=-\frac{i}{\hbar}\sum_{l}V_{kl}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)A_{l},$ (23)
We want to make two important observations on properties of Eq. (23). First, this equation describes unitary evolution, since on a time interval that tends to zero, we denote it as $dt$, the change of the column of amplitudes is given by
$\mathcal{A}(t+dt)=e^{\mathcal{M}(t)dt}\mathcal{A}(t),$
$M_{kl}(t)=-\frac{i}{\hbar}V_{kl}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right),$ (24)
which is unitary transformation since $M_{kl}=M_{lk}^{\dagger}$. On a finite time interval, transformation of amplitudes can be expressed as an infinite product of exponents of the form (24), so it is unitary as a product of unitary transforms.
Second, if a diagonal element of the interaction matrix is nonzero, $V_{kk}\neq 0$, which means self-interaction, we can make a substitution $A^{\prime}_{k}=A_{k}\exp(\frac{i}{\hbar}V_{kk}t)$, so that the diagonal term in Eq. (23) cancels out. Introducing exponential multipliers in such a form to the wave function (21) is equivalent to changing the energy of state $|k\rangle$ by $V_{kk}$. This coincides with the correction to energy in the first order of time-independent perturbation theory. Further, we will assume energies of states already include self-interaction and $V_{kk}=0$.
Equation (23) is too complicated to tackle directly. To continue our investigation, we need to use an approximate method. The original Diracs approach is to employ the variation of parameters method. We choose a different approach here.
Equation (23) is too complicated to tackle directly. To continue our investigation, we need to use an approximate method. The original Dirac’s approach is to employ the variation of parameters method. We choose a different approach here.
Let us assume that the amplitudes of the wave function are slowly changing over time compared to the fast oscillating exponent $\exp\left(i(E_{k}-E_{l})t/\hbar\right)$ in (23), for most pairs of values of $k$ and $l$. We want to distinguish the slowly oscillating terms of $\mathcal{M}(t)$ with $E_{k}\approx E_{l}$ that contribute to the change of amplitudes, from the fast oscillating terms that should affect only the fast oscillating part of the wave function. To do that, we integrate over some technical timescale $\Delta t$ that is small compared to the typical timescale of change of $A_{k}$, which we denote as $T_{A}$. However, $(E_{k}-E_{l})\Delta t/\hbar\gg 1$ for most values of $k$ and $l$.
Our general strategy for the subsequent derivation is to use the technical timescale $\Delta t$ as an intermediate value between the characteristic timescales of the fast-changing part of the wavefunction $|k\rangle$ and the slowly varying envelope $A_{k}$ (Fig. 2a). When integrating the microstate wavefunction $|k\rangle$ over $\Delta t$, we treat $\Delta t$ as a large quantity, almost infinity. However, when we deal with the amplitude $A_{k}$, we treat $\Delta t$ as a small quantity, nearly zero. Nevertheless, in both cases, we cannot simply take the limits $\Delta t\to 0$ or $\Delta t\to\infty$, because we would lose important information by doing so. We need to keep it finite and take the limits only at the very last steps of our derivation. In the end, we will see that if such an intermediate timescale exists, the results we obtain do not depend on the specific value of $\Delta t$.
After integration over $\Delta t$, on the left side of the equation (23), we obtain $A(\Delta t)-A(0)$. On the right side,
[…]
We divide both sides by $\Delta t$. We define the obtained expression before $A$ in the right side of the equation as $\chi_{\Delta E}(E_{k})$, where $\Delta E=|E_{k}-E_{l}|$. We will consider two cases:
$\Delta E\Delta t\gg\hbar=>\chi_{\Delta E}(E_{k})\approx 0,$
$\Delta E\Delta t\ll\hbar=>\chi_{\Delta E}(E_{k})\approx 1.$ (26)
We see that $\chi_{\Delta E}(E_{k})$ guarantees that only states with energies close to $E_{k}$ contribute to the variation of $A_{k}$. We want to emphasize this point even more and simplify our calculation. To do that, we replace $\chi_{\Delta E}(E_{k})$ with an indicator function $\overline{\chi}_{\Delta E}(E_{k})$, which equals one in a region of energies near $E_{k}$ and zero outside of this region.
[…]
$\Delta E\Delta t>$ $\pi\hbar=>\overline{\chi}_{\Delta E}(E_{k})=0$
$\Delta E\Delta t<$ $\pi\hbar=>\overline{\chi}_{\Delta E}(E_{k})=1.$ (27)
Later, we will see what information we have lost with this simplification.
On the left side of the equation (23), after performing transformations, we have
$\frac{A(\Delta t)-A(0)}{\Delta t}\approx\dot{A},$ (28)
but here we have a time derivative of smoothed $A$, which we have truncated the fast oscillating terms from.
We come to the equation for amplitudes, that are smooth, slowly oscillating part of the total wave function:
$\dot{A}_{k}=-\frac{i}{\hbar}\sum_{l}\chi_{\Delta E}(E_{k})V_{kl}A_{l}.$ (29)
Later we will not write $\chi_{\Delta E}(E_{k})$ explicitely, but will keep in mind that we sum over $l$-s such that $E_{k}-E_{l}<\Delta E$ (Fig. 2b). In the matrix form the equation is
$\dot{\mathcal{A}}=-\frac{i}{\hbar}\mathcal{V}\mathcal{A}$ (30)
This equation is much simpler than the one we started from in (23), because the matrix elements are now time-independent. However, we note that the matrix $\mathcal{V}$ is Hermitian, so equation (30) still describes unitary evolution of the system. Thus, we have not lost this fundamental property through our simplifications.
Using equation (30), we can estimate the characteristic timescale of the evolution of amplitudes, which we have denoted as $T_{A}$. We take the smallest nonzero element of the matrix $\mathcal{V}$ and denote it as $V_{M}$. Then the minimum possible rate of change of an amplitude is proportional to $V_{M}$, and we can write $T_{A}=\hbar/V_{M}$ (Fig. 2a).
Figure 2: The diagram illustrates the characteristic scales of transitions in a non-coherent regime. To make the diagram easier to understand, quantities that are an order of magnitude greater than others are shown as only slightly greater. Both the time scale and energy scale are depicted, as some proportions are more intuitive on the timescale, while others are more intuitive on the energy scale. The quantities on the time scale and energy scale are related by:
$T_{k}=\frac{\hbar}{E_{k}}$, $t_{\text{coh}}=\frac{1}{\omega_{\text{idth}}}$, $\Delta t=\frac{\pi\hbar}{\Delta E}$, $T_{A}=\frac{\hbar}{V_{M}}$.
(a) shows a microstate $|k\rangle$ and the probability amplitude of the system to be in such a microstate, $A$. It demonstrates that the characteristic scale of the amplitude evolution, $T_{A}$, is much higher than the working scale, $\Delta t$, which is much higher than the wave period of a state, $T_{k}$. Therefore, we have:
$T_{k}\ll\Delta t\ll T_{A}$.
(b) depicts a sheaf of states (represented as an oval), with a frequency width $\omega_{\text{idth}}$, containing a single state $|k\rangle$ in the first sheaf and a group of states in the second sheaf, indexed by $l$, such that the energy difference between them and the initial state is less than $\Delta E$. The width of the sheaf is much greater than the working broadening, $\Delta E$. It is evident that the number of states to which a transition can occur from the initial state $|k\rangle$ is proportional to the broadening $\Delta E$.
The solution to Eq. (30) is
$\mathcal{A}(t)=\exp\left(-\frac{i}{\hbar}\mathcal{V}t\right)\mathcal{A}(0)$ (31)
We may present the exponent as an infinite product and write
$\mathcal{A}(t)=\lim_{n\to\infty}\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}t/n\right)^{n}\mathcal{A}(0)$ (32)
The strategy of our derivation is to take $n$ as a large number, but not instantly take the limit $n\to\infty$. If we take this limit we would get $\dot{P}=0$. Instead, we want to expand functions to the Taylor series in the precision to first non-vanishing terms in $1/n$ and analyze its asymptotic behavior at $n\to\infty$. We denote $\Delta t=t/n$. We rewrite (32) as
$\mathcal{A}(t)=\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}\Delta t-\frac{1}{2\hbar^{2}}\mathcal{V}^{2}\Delta t^{2}+O(\Delta t^{3})\right)^{n}\mathcal{A}(0).$ (33)
Soon we will see the importace of keeping expression of order $\Delta t^{2}$.
We now want to analyze the non-coherent regime. This means that the state under consideration, $|k\rangle$, belongs to a sheaf of states with a width in the frequency region $\omega_{\text{idth}}$, and we choose a working timescale, $\Delta t$, such that it is much larger than the coherence time. Equivalently, we have $\Delta E\ll\omega_{\text{idth}}$ (Fig. 2b). The quantity $\Delta E$ fits into $\omega_{\text{idth}}$ many times, each time with a different phase difference for each pair of amplitudes, so we can average over the phase. This allows the formula for a series of non-coherent transformations, given in (20), to be applicable.
Let us write a product formula for $\mathcal{P}$
$\mathcal{P}(t)=\mathcal{T}^{n}\mathcal{P}(0),$ (34)
Similarly to what we have done with unitary matrices in expression (32), we want to express bistochastic matrices $\mathcal{T}$ in a form of unity matrix plus $\mathcal{Q}\Delta t$. Using (6) we have
$\mathcal{T}=\mathcal{E}+\mathcal{Q}\Delta t=|\mathcal{E}-\frac{i}{\hbar}\mathcal{V}\Delta t|^{\circ 2},$ (35)
For the non-diagonal terms of $\mathcal{Q}$ we find that
$k\neq l,\ \ \ Q_{kl}=\frac{1}{\hbar^{2}}|V_{kl}|^{2}\Delta t.$ (36)
For diagonal terms unit matrices cancel out and we have to calculate the diagonal of the matrix
$Q_{kk}=\left(\frac{\Delta t}{\hbar}\right)\left(\frac{1}{2}\mathcal{V}^{2}+\frac{1}{2}\mathcal{V}^{*^{2}}\right)_{kk}.$ (37)
For this we use the fact that $\mathcal{V}$ is Hermitian with zero diagonal elements, so we obtain
$Q_{kk}=-\frac{1}{\hbar^{2}}\left(\sum_{l}|V_{kl}|^{2}\right)\Delta t.$ (38)
We can see that this satisfies the conditions for the transition rate matrix of a CTMC. The non-diagonal elements are positive numbers, and the diagonal elements are negative, such that the sums of the elements in each column are zero. We observe that if we had not retained elements of order $\Delta t^{2}$ in (35), we would have zeros on the main diagonal of $\mathcal{Q}$.
We rewrite the formula (35) as $\exp(\mathcal{Q}t)(1+O(1/n))$. Now, recalling that $\Delta t=t/n$ is a small quantity compared to the characteristic timescale of change of $\mathcal{P}$, we have $1/n\ll 1$, and we can discard it. Thus, we obtain:
$\mathcal{P}(t)=\exp(\mathcal{Q}t)\mathcal{P}(0).$ (39)
Which means $P(t)$ is the solution of the differential equation
$\dot{\mathcal{P}}=\mathcal{Q}\mathcal{P}.$ (40)
This is the equation of non-coherent evolution that we were looking for.
We can reason differently. Instead of using the solution for $\mathcal{A}(t)$ to construct $\mathcal{P}(t)$ and then deducing the equation for $\mathcal{P}(t)$, we aim to derive the differential equation directly. Since the probability of the system being in state $k$ is $P_{k}=A_{k}^{*}A_{k}$, its time derivative is normally given by $\dot{P}_{k}=\dot{A}_{k}^{*}A_{k}+A_{k}^{*}\dot{A}_{k}$. However, we use the smoothed derivative of the amplitude, as given by (28), which is not defined at a single point but uses two different points in time. This introduces an ambiguity: we can substitute $A_{k}(0)$ or $A_{k}(\Delta t)$ into the formula, or their linear combination, or even take the value of the amplitude at any point within the interval between 0 and $\Delta t$, and this yields different answers for $\dot{P}_{k}$. To overcome this difficulty, we use the smoothed derivative of the probability, just as we used the smoothed derivative of the amplitude in (28). We regroup the terms so that the formula resembles the normal one based on the product rule, while also being symmetric with respect to the transposition of $A_{k}^{*}$ and $A_{k}$. We omit the $k$ index and write
$\dot{P}=\frac{P(\Delta t)-P(0)}{\Delta t}=\frac{A^{*}(\Delta t)A(\Delta t)-A^{*}(0)A(0)}{\Delta t}=$
$=\frac{[A^{*}(\Delta t)-A^{*}(0)]A(\Delta t)/2+A^{*}(0)A(\Delta t)/2+A^{*}(\Delta t)[A(\Delta t)-A(0)]/2+A^{*}(\Delta t)A(0)/2-A^{*}(0)A(0)}{\Delta t}=$
$=\dot{A}^{*}\frac{A(\Delta t)+A(0)}{2}+\frac{A^{*}(\Delta t)+A^{*}(0)}{2}\dot{A}.$ (41)
We note that the first and the last expressions are equivalent, both can be used for calculation. We introduce the last form because it resembles the product rule formula. And we assume we use the last form in subsequent manipulations.
For $\mathcal{A}(\Delta t)$ we use the formula (35), setting $t=\Delta t$
$\mathcal{A}(\Delta t)=\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}\Delta t-\frac{1}{2\hbar^{2}}\mathcal{V}^{2}\Delta t^{2}+o(\Delta t^{2})\right)\mathcal{A}(0).$ (42)
This time we want to write up the computation element-wise, to present it both ways. For $A_{k}(\Delta t)$, only keeping terms of the first order in $\Delta t$:
$A_{k}(\Delta t)=A_{k}(0)-\frac{i}{\hbar}\sum_{l}V_{kl}\Delta tA_{l}(0)$ (43)
Then for derivative, to the first non-vanishing term in $\Delta t$
$\dot{A}_{k}(\Delta t)=-\frac{i}{\hbar}\sum_{l}V_{kl}A_{l}(0)+\frac{1}{\hbar^{2}}\left(\sum_{l}|V_{kl}|^{2}\right)\Delta tA_{k}(0).$ (44)
We can write equivalent expressions for $A^{*}_{k}$ and its derivative.
We substitute expressions (43, 44) and their complex conjugates into (41) and imply non-coherence condition that makes $A^{*}_{k}A_{l}\rightarrow\delta_{kl}|A_{k}|^{2}=P_{k}$. Thus we obtain
$\dot{P}_{k}=-\frac{1}{\hbar^{2}}\left(\sum_{l}|V_{kl}|^{2}\right)\Delta tP_{k}+\frac{1}{\hbar^{2}}\sum_{l}|V_{kl}|^{2}\Delta tP_{l}.$ (45)
Taking into consideration the definition of the $\mathcal{Q}$ matrix (36, 38), we see that it is the same equation as (40). We observe that we obtain the same result with a slightly different reasoning. In the latter case, the connections with the results of the previous section are less apparent, but the differential equation arises more naturally.
Let us now address the issue with formula (40) (or its equivalent (45)): its dependence on the arbitrarily chosen parameter $\Delta t$. We will demonstrate that, with the chosen approximations, the system’s dynamics do not depend on the choice of $\Delta t$.
We begin by utilizing an approximation of a continuous spectrum. Near a given state $l$, we divide the quantum numbers into two components. The first is $l_{T}$, such that small changes in its value leave the energy unchanged, so it is tangential to the energy surface. The second is $l_{N}$, where a change results in a change in energy. We express energy as a function of the quantum numbers in the form $E(l)=E(l_{0})+l_{N},\mathrm{grad}_{l}E$. Additionally, we recall expression (27) and substitute $\Delta t=\pi\hbar/\Delta E$. We then obtain:
$\frac{1}{\hbar^{2}}\sum_{l}|V_{kl}|^{2}\Delta tP_{l}=\frac{\pi}{\hbar\Delta E}\sum_{l_{T},l_{N}}|V_{kl}|^{2}P_{l_{T},l_{N}}$ (46)
If $P_{l_{N}}$ is a smooth function of energy at the point of interest, in a short interval we can approximate it as a constant, so summation over $l_{N}$ gives the number of states inside the energy region – $N$ (Fig. 2b):
$\frac{\pi}{\hbar\Delta E}\sum_{l_{T},l_{N}}|V_{kl}|^{2}P_{l_{T}}=\frac{\pi N}{\hbar\Delta E}\sum_{l_{T}}|V_{kl}|^{2}P_{l_{T}}=\frac{2\pi}{\hbar}\sum_{l_{T}}|V_{kl}|^{2}P_{l_{T}}\nu(E)$ (47)
where $\nu(E)=N/(2\Delta E)$ is the dencity of states. Multiplier 2 arises because $\Delta E$ is the absolute value of the deviation of energy from $E_{l}$ and the energy in the region can be less than $E_{l}$ and higher than $E_{l}$ so the length of the region is $2\Delta E$.
We can reason differently. We perform the same subsitution $\Delta t=\pi\hbar/\Delta E$ as in (46), and approximate the sum as an integral:
[…]
where $C$ is some constant with dimension $[1/l]$ that arises due to the transition from the sum to the integral. We observe that $\frac{\overline{\chi_{\Delta E}}}{2\Delta E}$ is a function that equals $1/2\Delta E$ in the energy region near $E_{K}$ with width $\Delta E$, and is zero elsewhere. As $\Delta E\to 0$, the region of energies shrinks to zero, but the value of the transmission rate grows to infinity. We can see that the integral of $\frac{\overline{\chi_{\Delta E}}}{2\Delta E}$ is a function that equals $1/2\Delta E$ over energy, and its total integral is one. Thus, in the weak limit, the function approaches $\delta(E-E_{k})$. Therefore, for transmissions between neighborhoods of states $k$ and $l$, we obtain:
$Q_{kl}=\frac{2\pi}{\hbar}|V_{kl}|^{2}\delta(E_{k}-E_{l}).$ (49)
That finishes our derivation.
Here we can see what we have lost by using simplification (27): the limit of $\frac{\chi_{\Delta E}}{2\Delta E}$ at $\Delta E\to 0$ is not just a delta function, but a sum $\delta(E-E_{k})+\text{v.p.}\frac{i}{E-E_{k}}$, where v.p. means that taking an integral over $E$ one should take its principal value. This expression is very well-known in the context of perturbation theory. The second term is responsible for perturbational corrections to the values of energies of states and matrix elements of interaction.
We could have chosen a more rigorous method for deriving equation (30). We would still seek a smoothed unitary evolution equation, which we achieve by expressing the matrix $\mathcal{V}$ in the form of the Dyson series. If we only retain the first term of the series for $\mathcal{V}$, it gives the same expression that we have found. Seemingly, this method allows us to find higher-order terms of the perturbation series, including corrections to energy levels due to interactions and second-order transitions. However, this approach is lengthy, and if we were to take this route, we would aim to investigate perturbation theory within the described framework. In this paper, our primary goal is to deduce the non-coherent evolution equation from the smoothed unitary evolution equation (30). We leave the task of finding the most accurate form of the smoothed unitary evolution equation, which includes perturbative corrections, for subsequent publications.
Our analysis began with partitioning the wave function of the system into two components: fast-varying time-dependent wavefunctions of microstates and slow-varying probability amplitudes for the system to be in a given state, as shown in (21) (Fig. 2a). It is important to note that the distinction between coherent and non-coherent evolution regimes can only be made at the level of smooth amplitudes. The microstates we consider are combinations, either symmetric or anti-symmetric, of the exact stationary solutions to the single-particle Schrödinger equation. At this level, we cannot even introduce the concepts of coherence or non-coherence. The situation changes when we consider envelopes that contain a sheaf of states within a small energy interval and the interactions between them (Fig. 2b). As Eq. (11) shows, sheaves can be phase-correlated, which gives the regime of unitary evolution (30), or completely uncorrelated, which leads to the regime of non-coherent evolution (40).
We see that if we do not consider the full set of equations (40) describing the evolution of the system’s state through transitions between microstates, but instead focus on the single transition rate (49), we recover the well-known Fermi rule. It is worth noting that the arguments we have presented apply not only to macroscopic closed systems but also, for example, to a single atom emitting a photon into space. The Fermi rule was originally derived for such a situation.
Indeed, we have made two crucial assumptions in our derivation. The first is the applicability of averaging over phases. If the atom and the object with which the emitted photon would interact are not phase-correlated, the second non-coherence mechanism described in the previous section (Fig. 1c) comes into play, allowing averaging to be applicable. The second assumption is the continuous spectrum. While an atom has a discrete spectrum, the system consisting of the atom and the electromagnetic field has a continuous spectrum, so the arguments hold. The key difference between an isolated system and an atom emitting into empty space is that once the photon departs from the atom, no other photons follow, thus disrupting the symmetry of forward and backward transitions.
Since the Bohr model of the atom, the notion of ”quantum jumps” has been present in quantum theory. This is an ironic term for transitions between states of the system that, according to the Bohr model, occurred instantaneously. The theory of decoherence describes ”quantum jumps” as a continuous process of decoherence happening locally on a very short timescale [44]. Our description of quantum transitions allows us to estimate their characteristic timescales.
We have derived the expression (40), where transitions between states are assumed to be instantaneous. However, our derivation allows us to see that the transition time is only negligibly small compared to the characteristic timescale of the evolution of amplitudes. Let us consider the emission of a photon. More generally, we can discuss a transition between fermionic states with the emission of a wave — a boson.
First, we observe that starting from the initial state $|k\rangle$, we can transition equally to two different states: one containing a boson with energy smaller than $\Delta E$, and the other not (Fig. 2b). Thus, such states are ”glued” in our formalism. Hence, bosons with energies less than $\Delta E$ are practically indistinguishable from the absence of any bosons. On a timescale $\Delta t$, we can only ”detect” bosons with energy $\hbar\omega\geq\hbar/\Delta t$, meaning $T\leq\Delta t$, where $T$ is the wave period. On the other hand, we cannot separate events in time that are separated by less than $\Delta t$. The transition time is the minimal time interval that still allows the detection of the emission: $\tau=\mathrm{min}{\Delta t:T\leq\Delta t}$. This implies $\tau=T$; the transition time is equal to the period of the emitted wave, which makes perfect sense.
In the next section, we will suggest a way to include the waves of small frequency within the presented framework.
IV Applications to statistical mechanics
In the previous section, we deduced the equation of stochastic evolution and, by considering a single element of the transition matrix, derived the Fermi rule. We could reason in the opposite direction, as was done in Ref. [14]. If we combine all the transition rates into a matrix and add the diagonal elements to satisfy the probability conservation condition, we obtain a stochastic equation known as the Fermi master equation. It coincides with the non-coherent evolution equation (45). This provides a possible way to deduce Eq. (45) without reference to non-coherence. However, it lacks proper justification.
We want to use the derived equation to address the problem of irreversibility in statistical mechanics. We consider all three instances of this problem mentioned in the introduction: the second law of thermodynamics, the equidistribution of states in equilibrium, and the irreversibility of the Boltzmann equation. In [14], it is proven that entropy, defined as
$S=-\sum_{k}P_{k}\ln P_{k}$ (50)
increases in a system whose evolution is described by the Fermi master equation, unless it has already reached its maximum. Entropy, defined by equation (50), attains its maximum when the probabilities of all states are equal. Thus, this also proves the equiprobability of states in equilibrium. However, we can approach the problem from the opposite direction.
As we mentioned earlier, the equation of non-coherent evolution (40) is a continuous-time Markov chain (CTMC) equation. It is known [42] that such equations are irreversible and converge to a stationary distribution for large values of $t$. The specific distribution depends on the particular $\mathcal{Q}$-matrix. However, the $\mathcal{Q}$-matrix of non-coherent evolution, defined by expressions (36, 38), is symmetric. For symmetric $\mathcal{Q}$-matrices, the stationary distribution is always the same: an equiprobability of all microstates.
The symmetry of the $\mathcal{Q}$-matrix aligns with the principle of detailed balance. This symmetry is a consequence of the Hermiticity of the interaction matrix $\mathcal{V}$.
In contrast with the general case proven in [42], the symmetric case is simpler, and we can directly verify that the equiprobability of all microstates is a stationary solution of equation (40). We substitute $P_{k}=1/N$ for each $k$ into the system (40), where $N$ is the number of equations (i.e., the number of microstates). Since the sum of the elements in each row of the symmetric $\mathcal{Q}$-matrix equals zero, the right-hand side of (40) vanishes. Therefore, the time derivatives of all $P_{\alpha}$ equal zero. This shows that the distribution $P_{\alpha}=1/N$ is stationary, and regardless of the initial condition, in the limit $t\to\infty$, the probability of each state becomes equal. This means the distribution to which the non-coherent evolution of a closed macroscopic system converges is the Gibbs microcanonical ensemble.
We note that entropy growth is a stronger property of non-coherent evolution than irreversibility alone. All CTMCs are irreversible; however, only symmetric ones have the equiprobability of states as a stationary solution. Moreover, equiprobability corresponds to the maximum of entropy as defined by equation (50). Thus, entropy growth is a consequence of irreversibility combined with the principle of detailed balance.
We also emphasize that the approximation (27) neglects the perturbation of states due to interaction. We may speculate that even when interactions are taken into account, the system still undergoes stochastic evolution described by equation (40), but in terms of microstates with renormalized energies and interaction matrix elements. This conjecture is motivated by the fact that Gibbs’ theory remains valid for systems with strong interactions. Since we have shown that equation (40) leads to the microcanonical ensemble, it is plausible that for general systems the equidistribution of microstate probabilities results from the evolution equation being a symmetric CTMC. However, we are not in a position to prove this rigorously, and therefore we proceed under the assumption of weak interaction, for which approximation (27) remains valid.
Now we want to derive the Boltzmann collision integral. In the second section, we have described two mechanisms leading to the lack of coherence in interaction. Here we consider two particular types of interactions: the emission and absorption of phonons by electrons and the three phonons interaction. In the third section we found that for our theory to be applicable, the mean time of quantum transitions should be much higher than the coherence time: $t_{\text{coh}}/T_{A}\ll 1$. If we switch from times to lengths, the mean free passes of particles are much higher than the coherence lengths. If also the size of the system is much higher than the mean free paths of the particles, most interactions in the systems occur between particles whose emissions are not causally related. In the first example, a phonon is typically absorbed by a different electron than the one that emitted it. In the second example, the interaction of phonons occurs between phonons that were emitted at a distance much greater than the coherence length from each other. It gives us right to perform the averaging over phases thus leading to the applicability of the non-coherent evolution formula (40).
Let us consider the first case: the transitions between energy levels by electrons with the emission or absorption of phonons. More generally, it could be any fermions and bosons. The interaction has the form $V_{ij}\hat{c}^{\dagger}_{i}\hat{c}_{j}\hat{a}$, where $\hat{c},\hat{a}$ are annihilation operators for fermions and bosons, respectively, and indices $i,j$ denote fermionic single-particle states.
Let us first consider just two fermionic states and the transitions between them. We denote states of a noncoherent system by angular brackets on both sides; in this notation, we consider transitions $\langle 0,1,n-1\rangle\leftrightarrow\langle 1,0,n\rangle$. Numbers in the brackets are the occupation numbers of fermionic states 1 and 2, respectively, and the third one is the number of bosons. We use formula (45) with $k=\langle 0,1,n-1\rangle$, $l=\langle 1,0,n\rangle$, and $q=2\pi|V_{1,2}|^{2}\nu/\hbar$:
$\dot{P}_{\langle 0,1,n-1\rangle}=-nqP_{\langle 0,1,n-1\rangle}+nqP_{\langle 1,0,n\rangle}$
$\dot{P}_{\langle 1,0,n\rangle}=-nqP_{\langle 1,0,n\rangle}+nqP_{\langle 0,1,n-1\rangle}.$ (51)
We want to find how the mean occupation number (which is a one-particle distribution function) of the first state $f_{1}$ varies in time. We write down the expression $f_{1}$ through probabilities
$f_{1}=$ $P_{\langle 0,1,0\rangle}\times 0+P_{\langle 0,1,1\rangle}\times 0+P_{\langle 0,1,2\rangle}\times 0+...$
$+P_{\langle 1,0,0\rangle}\times 1+$ $P_{\langle 1,0,1\rangle}\times 1+P_{\langle 1,0,2\rangle}\times 1+...$ (52)
Vertically aligned are states between which the transitions occur.
We take time derivative from both sides and using (51) we get
$\dot{f}_{1}=(-qP_{\langle 1,0,1\rangle}+qP_{\langle 0,1,0\rangle})+(-2qP_{\langle 1,0,2\rangle}+2qP_{\langle 0,1,1\rangle})+...=\sum_{n}nq(P_{\langle 0,1,n-1\rangle}-P_{\langle 1,0,n\rangle}).$ (53)
To deduce the one-particle distribution function from many-particle probabilities, we need to make a simplifying assumption. We assume that the probabilities of different occupation numbers of different states are independent of each other. That means $P_{\langle l,m,n\rangle}=P_{\langle l,...\rangle}P_{\langle...,m\rangle}P_{\langle n\rangle}$. For the mean occupation number of phonons $N$ we have
$N=\sum_{n=0}nP_{\langle n\rangle}$ (54)
We substitute this expressions into (55) and obtain:
$\dot{f}_{1}=\sum_{n}nq(P_{\langle 0,...\rangle}P_{\langle...,1\rangle}P_{\langle n-1\rangle}-P_{\langle 1,...\rangle}P_{\langle...,0\rangle}P_{\langle n\rangle})=q\sum_{n}n\left[(1-f_{1})f_{2}P_{\langle n-1\rangle}-f_{1}(1-f_{2})P_{\langle n\rangle})\right]=$
$=q\left[(1-f_{1})f_{2}\sum_{k=0}(k+1)P_{\langle k\rangle}-f_{1}(1-f_{2})\sum_{k=0}kP_{\langle k\rangle}\right]=q\left[(1-f_{1})f_{2}(N+1)-f_{1}(1-f_{2})N\right].$ (55)
Summation over all possible transitions yields the Boltzmann collision integral [45].
Let us derive the collision integral for the case of the three-phonon process. We consider three bosonic states, with occupation numbers $n_{1},n_{2},n_{3}$. The sum of the energies of the first two states is equal to the energy of the third state, and interaction permits bosons of the first two states to merge into a boson in a third state and vice versa. So transitions under consideration are $\langle n_{1}-1,n_{2}-1,n_{3}+1\rangle\leftrightarrow\langle n_{1},n_{2},n_{3}\rangle\leftrightarrow\langle n_{1}+1,n_{2}+1,n_{3}-1\rangle$. With formula (45) we have:
$\dot{P}_{\langle n_{1},n_{2},n_{3}\rangle}=qn_{1}n_{2}(n_{3}+1)P_{\langle n_{1}-1,n_{2}-1,n_{3}+1\rangle}+q(n_{1}+1)(n_{2}+1)n_{3}P_{\langle n_{1}+1,n_{2}+1,n_{3}-1\rangle}-$
$q\left(n_{1}n_{2}(n_{3}+1)+(n_{1}+1)(n_{2}+1)n_{3}\right)P_{\langle n_{1},n_{2},n_{3}\rangle}.$ (56)
We take the time derivative of Eq. (54) and substitute Eq. (56) into it to get
$\dot{N}_{1}=q\sum_{n_{1},n_{2},n_{3}}n_{1}[n_{1}n_{2}(n_{3}+1)P_{\langle n_{1}-1,n_{2}-1,n_{3}+1\rangle}+(n_{1}+1)(n_{2}+1)n_{3}P_{\langle n_{1}+1,n_{2}+1,n_{3}-1\rangle}-$
$\left(n_{1}n_{2}(n_{3}+1)+(n_{1}+1)(n_{2}+1)n_{3}\right)P_{\langle n_{1},n_{2},n_{3}\rangle}]$ (57)
We make independence assumption and sum over $n_{2},n_{3}$ using formula (54) and shifting indices similarly to what we did with expression (55), to get
$\dot{N}_{1}=q\sum_{n_{1}}n_{1}[n_{1}P_{\langle n_{1}-1\rangle}(N_{2}+1)N_{3}+(n_{1}+1)P_{\langle n_{1}+1\rangle}N_{2}(N_{3}+1)-$
$n_{1}P_{\langle n_{1}\rangle}N_{2}(N_{3}+1)-(n_{1}+1)P_{\langle n_{1}\rangle}(N_{2}+1)N_{3}]$ (58)
We regroup summands, take $n_{1}$ into brackets, shift indices and obtain
$\dot{N}_{1}=q\sum_{n_{1}}\left((N_{2}+1)N_{3}[(n_{1}+1)^{2}-n_{1}(n_{1}+1)]+N_{2}(N_{3}+1)[(n_{1}-1)n_{1}-n_{1}^{2}]\right)P_{\langle n_{1}\rangle}$ (59)
Summation over $n_{1}$ yeilds
$\dot{N}_{1}=q(N_{1}+1)(N_{2}+1)N_{3}-qN_{1}N_{2}(N_{3}+1)$ (60)
Summation over all possible combinations of the second and third states gives the first part of the collision integral for the three-phonon processes. The second part is obtained similarly if we consider the decay of phonons in state one into two phonons of states two and three and vice versa. It gives contribution of the form
$q(N_{1}+1)N_{2}N_{3}-qN_{1}(N_{2}+1)(N_{3}+1)$. The calculation is completely analogous [45].
Figure 3: The diagram shows how the Boltzmann equation and the Gibbs microcanonical ensemble are deduced from the equation of non-coherent evolution (Eq. 45). In rectangles, the names of equations are indicated, and arrows represent the mathematical procedures used to derive a formula from a more general one. $t\to\infty$ denotes taking the stationary point of the equation, and $\mathbb{P}\to\mathbb{E}$ indicates averaging – that is, the transition from probabilities to average (expected) values. The diagram is commutative in the sense that taking the stationary point of Eq. (45) and then averaging, or averaging first and then taking the stationary point, both lead to well-known single-particle distribution functions (Fermi-Dirac or Bose-Einstein distributions).
Thus, we have derived two cases of the Boltzmann collision integrals. With this derivation, the irreversibility of the Boltzmann equations comes as no surprise, since the irreversibility of the collision part is inherited from the irreversibility of the non-coherent evolution equation (40), from which it is derived.
During the derivation of the non-coherent evolution equation (40), we noted that we should pick $\Delta t$ small compared to the typical timescale of the change of probability amplitude: $\Delta t\ll T_{A}$ (Fig. 2). At the end of the previous section, we observed that the period of emitted waves in a given transition, $T_{ij}$, should be less than $\Delta t$, which yields $T_{ij}\ll T_{A}=\hbar/V_{ij}$. We can see that for long waves with period $T_{ij}\to\infty$, this condition cannot be fulfilled unless the matrix element of interaction with long-wavelength bosons, $V_{ij}$, tends to zero as the wavelength goes to zero. Additionally, in the previous paragraph, we noted that low-frequency waves were excluded from our derivations. We suggest that low-frequency waves contribute to the force term in the Boltzmann equation and, therefore, should be treated separately. In this paper, we have derived the interaction in a local region, which corresponds to the collision term in the Boltzmann equation. The free movement of particles in the media, corresponding to the diffusion term, and interaction with low-frequency fields, corresponding to the force term, should be derived separately. One possibility is the approach described in Ref. [7].
In this section, we have derived the Gibbs microcanonical ensemble, which forms the basis for all equilibrium statistical mechanics, but exclusively in the case of weakly interacting systems, or simply gases. We also derived the collision part of the Boltzmann equation, which is the fundamental equation of the physical kinetics of gases. It is interesting to note that the one-particle distribution functions of quantum gases (Fermi-Dirac or Bose-Einstein distributions) can be derived in two ways: by finding the mean occupation number from the microcanonical ensemble or by finding the stationary point of the Boltzmann collision integral. In the present paper, we have derived the Gibbs microcanonical distribution as a stationary point of the equation of non-coherent evolution (Eq. 45), and the Boltzmann collision integral as the result of the transition from the evolution equation for probabilities to the equation for average values. This means that if we start from Eq. 45, we can first take the stationary point and then perform the averaging, or first take the averages and then find the stationary point. Both approaches lead to the same one-particle distribution functions. This result is depicted in the diagram in Fig. 3.
We want to add that typically, the canonical ensemble is derived from the microcanonical ensemble, and then everything in equilibrium statistical mechanics is derived from the canonical ensemble. However, the one-particle distribution functions can be derived directly from the microcanonical ensemble (see Ref. [5]). In the canonical ensemble, the probabilities of different occupation numbers of different states are inherently independent of each other. During the derivation from the microcanonical ensemble, this assumption is introduced, in the same way we did in our derivation of the Boltzmann collision integral.
V Summary and conclusions
The present work develops a theoretical framework connecting quantum mechanics with irreversible statistical processes through the concept of non-coherent evolution. We have demonstrated how averaging over phases in quantum systems leads to the emergence of irreversible behavior, providing a bridge between reversible quantum mechanics and irreversible statistical mechanics. We also provided an explanation of how this averaging arises.
We can loosely divide the manuscript into two parts. In the first part (Section 2 and most of Section 3), we derive the equation of non-coherent evolution. In the second part (the end of Section 3 and Section 4), we use it to resolve problems associated with our understanding of quantum transitions and the irreversible behaviour of macroscopic systems.
At the beginning, we considered non-coherent transformations, showing how they naturally arise from unitary transformations when phase information is lost. Non-coherence arises from the fact that no wave has exactly one frequency but is composed of a sheaf of waves with closely spaced frequencies. Such sheaves can be mutually coherent or non-coherent. We identified two fundamental mechanisms leading to the loss of coherence between sheaves: temporal separation exceeding the coherence time and interaction between causally unrelated emissions. These mechanisms were illustrated using optical systems, where direct experimental verification is possible.
Then we move on from studying a discrete series of transformations to continuous evolution. We derive the non-coherent evolution equation, which is the central result of this work ((45), (40)). Our analysis begins with partitioning the wave function of the system into two components: fast-varying psi functions of microstates and probability amplitudes, which are smooth (in-time) envelopes of the psi function (21). Our derivation relies on the assumption of weak interaction and the existence of well-separated time scales. We introduced a technical timescale $\Delta t$, chosen between the microstate oscillation period and the characteristic scale of amplitude variation. Using this technical assumption, we showed that for timescales much larger than the coherence time, unitary evolution transforms into symmetric stochastic evolution, described by continuous-time Markov chains. We then show that the final results of the calculations do not depend on the value of the timescale $\Delta t$.
Importantly, our derivation reveals the connection between non-coherent evolution and Fermi’s golden rule, showing how the latter emerges naturally from our more general framework. Fermi’s rule calculates constant transition rates. Constant transition rates are a property of a stochastic process. For unitary evolution, in contrast, transition rates depend on the current values of amplitudes, as can be seen from Eq. (30). The unitarity of evolution, presented in the starting equation (23), is implicitly broken during the classical derivation of Fermi’s rule. With our methods, we explicitly identify the juncture in our derivation where the unitarity of the equations is broken.
There is also one dubious point in the classical derivation of Fermi’s rule: we integrate over an infinite time period to obtain a momentary transition rate. The introduction of the technical timescale makes this point very clear. This timescale is very large – ”infinite” – compared to the microstate oscillation period, and very small – ”infinitesimal” – compared to the characteristic scale of amplitude variation. Thus, we can treat it accordingly at different points in the derivation. Our approach also permits us to estimate the time interval of a quantum transition. It turns out that this time coincides with the period of the wave that is being emitted. Thus, we believe that the presented theory deepens our understanding of quantum transitions.
We continue by showing that the equation of non-coherent evolution serves as a common basis for both equilibrium statistical mechanics and kinetic theory. We derived the Gibbs microcanonical ensemble as its stationary solution, under the assumption of weak interaction, and showed that this stationary distribution corresponds to the equiprobability of microstates. We also derived the Boltzmann collision integrals by transitioning from equations for probabilities to equations for average values. This derivation justifies the irreversibility of the Boltzmann equation as a direct consequence of the irreversibility of non-coherent evolution. We note that the Boltzmann collision integral is derived strictly from the non-coherent evolution equation and basic definitions, without any ”hand-waving”.
A notable property of this derivation is the commutativity of taking average values and taking the stationary point of the evolution. If we perform these operations, we arrive at the single-particle distribution functions (Fermi-Dirac or Bose-Einstein distributions), and it does not matter in which order the operations are carried out. This highlights a deeper connection between the methods of equilibrium statistical mechanics and physical kinetics.
Our results suggest that the physical concept of non-coherence and the mathematical structure of symmetric continuous-time Markov chains are fundamental to understanding the emergence of irreversibility in quantum systems, potentially extending beyond the weak-interaction regime considered in the present manuscript.
Acknowledgements.
The author would like to thank Miguel Larotonda, who patiently answered my questions about coherence in the context of classical and quantum optics, and whose insightful responses greatly improved my understanding of these areas.